obtain the regression line (y=a+bx) from the following data

x: 5 9 1 10 5

y:7 10 3 8 2


1
Expert's answer
2021-08-02T12:59:50-0400

"\overline{x}=\frac{30}{5}=6,\:\:\overline{y}=\frac{30}{5}=6"

"\overline{x}=6\:and\:\overline{y}=6"


"b_{yx}=\frac{n\sum xy-\sum x\sum y}{n\sum x^2-\left(\sum x^2\right)}=\frac{5\left(218\right)-\left(30\times 30\right)}{5\left(232\right)-232}"


"\:b_{yx}=\frac{1090-900}{1160-232}=\frac{190}{928}=\frac{95}{464}"


Regression line "y" on "x"

"\:y-\overline{y}=b_{yx}\left(x-\overline{x}\right)"


"y-6=\frac{95}{464}\left(x-6\right)"


"464y-2784=95x-570"


"464y=95x+2214"


"464y=95x+2214"


"\:\:y=\frac{2214}{464}+\frac{95}{464}x"